OpenStudy (anonymous):

Log bottom thing 8 (.25)

OpenStudy (anonymous):

\[log_80.25\]

OpenStudy (anonymous):

that's just asking \[8^x=\frac{1}{4}\]

OpenStudy (anonymous):

okay thank you!!!!

OpenStudy (anonymous):

can you find x?

OpenStudy (curry):

x=-2/3